OpenStudy (anonymous):

Rewrite with only sin x and cos x sin 3x - cos x HELP!

OpenStudy (muhammad_nauman_umair):

elaborate it

OpenStudy (anonymous):

2 sin x - sin^3 x - cos x 2 sin x cos^2 x + sin x - 2 sin^3 x - cos x 2 sin^3 x cos^4 x + 1 3 sin x cos x - sin^3 x - cos x HERES THE ANSWER CHOICES (((((((:

OpenStudy (anonymous):

@Muhammad_Nauman_Umair

OpenStudy (anonymous):

@Zarkon help me please!

OpenStudy (anonymous):

or @saifoo.khan (:

OpenStudy (saifoo.khan):

sin(2x) = 2sinxcosx Now im stuck on how to use this for the left term in your problem.

OpenStudy (anonymous):

me too :(

OpenStudy (muhammad_nauman_umair):

either it is b or d option

OpenStudy (anonymous):

it was b thanks though

OpenStudy (saifoo.khan):

sin(3x) = sin(2x+x) = sin(2x) + sin(x)

OpenStudy (raden):

sin(a+b) = sinacosb +cosasinb

OpenStudy (saifoo.khan):

^

OpenStudy (muhammad_nauman_umair):

option b is true

OpenStudy (muhammad_nauman_umair):

As =Sin(2x+x)-Cosx =Sin2xCosx+Cos2xSinx-Cosx =(2 SinxCosx)Cosx+(1-2 Sin^2x)Sinx-Cosx =2SinxCos^2x+Sinx-2Sin^3x-Cosx

OpenStudy (anonymous):



OpenStudy (anonymous):

I got the answer guys don't worry! but thanks a ton(:

OpenStudy (anonymous):